# Investigation about a statement equivalent to Riemann Hypothesis (RH) applied to Dirichlet primitive L functions

### v15

@@ Appendix A Phase variations of $\xi(s)$, $\xi(s,\chi)$, $\zeta(s)$ and $L(s,\chi)$
With $\alpha_{1}=\alpha=0$, even $\chi_{primitive}$, or $\alpha_{1}=\alpha=1$, odd $\chi_{primitive}$. See (1.2 ). If we consider the factor (s1)(s-1) attached to $\zeta(s)$ like in (A.1), we can use (A.4) again putting $\alpha_{1}=0\ ,\ \alpha=2\ ,\ q=1$. As $z+1=\frac{s+\alpha}{2}\rightarrow z=\frac{2\epsilon+2\alpha-3}{4}+i\ \frac{t}{2}$, Stirling formula ( [7, p. 109,112]) allows us to write (A.4) for each of the 3 choices of $(\alpha,\alpha_{1})$:
@@
The $B_{j}$ are the Bernoulli numbers ( [7, p. 11] ), that vanish for odd $j$ while:
@@
where $\arg(z)$ is taken in the interval: $\pi<\arg(z)<\pi$ (see [7, p. 112].
@@
It is apparent that the difference between computation (3.13 ), upon substitution of (3.14) with (3.15) is absolutely convergent for sure from $\epsilon>0$ and conditionally convergent for $\epsilon=0$.
Proof. Let us look at (3.27), and, consider only one $h$-class primes till $p_{max}$ ( as in (3.18 ).
@@
If we consider all the $\phi(q)$ terms ( possible if $p_{max}<\infty$), even if we cannot conclude that the sum of all $\phi(q)$ contributions tend to zero ( as in Lemma 4), we can presume that the poor convergence of a single $h$-class primes is enanced when all $h$-classes are combined together. This seems confirmed by figg. 17, 18, and, 19.
@@ Appendix D Some feedback from readers
    ANSW.I agree on the necessity to fill this gap (not a mistake I think) in the logical path of the paper. In new version of section 3 I have shown that Euler product makes sense also in critical strip for primitive $L$-Functions. Equation (1.1) does not hold for $0<\Re(s)<1$, but, subsection 3.1 shows a way in order to exploit Euler product for variations, in particular phase variations, on small $t$-intervals. On the other hand ( even if it is not an argument) it seems to me worth to draw the attention on figg, 3 through 13 where it is apparent that (3.9), at least numerically, works rather well even with only $m=3$. In conclusion it seems to me that subsection 3.1 is a valid base for the developments of section 3. Anyhow I thank the reader for this observation that prompted me to delve deeper on subject.
@@
    ANSW. I agree with the statement, but it is not applicable.The “manipulation” of terms in (3.34 ) (or in others occasions (3.1), (3.18) and (3.20) ) is applied only at addends with index less than the running index. So on a set of finite numbers. Partial sums do not change (as it is stressed thoughout). To explain better. Consider an example $S_{j}=-1+1/2-1/3+1/4-.....=\sum_{j}(-1)^{j}/(j)\rightarrow-\ln(2)\ ,$ as $j\rightarrow\infty$.
    If we take the sum till $J_{max}$ and then we partition the addends with $j\leq J_{max}$ in positive $\sum^{+}_{J_{max}}$ and negatives $\sum^{-}_{J_{max}}$, then we can write $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$, and, we do not alter final result $\ln(2)$ for $J_{max}\rightarrow\infty$ transforming the limit in something like $\infty$-$\infty$. It is exactly the “manipulation” in expression : (3.34 ). While in (3.1), in (3.18), and, in (3.20), we manipulate to highlight some structure of the series, but , with indexes less than the running index . A completely rearranging of terms as in [1, p. 411] without change in final result is possible only where convergence is absolute. But, it is easy to see that there are infinite rearrangement of terms that preserve the sum $\ln(2)$ of example above. For example if we change odd with even terms, i.e. $S^{\prime}_{j}=-1/2+1-1/4+1/3-1/6+1/5-1/8+1/7-.....=-\ln(2)$.
@@
  • The argument in equations ( 3.43 ) through (3.50) rests on the idea that because each term in the series for $\epsilon>0$ is smaller in magnitude than the corresponding term for $\epsilon=0$ (due to the factor of $p^{-\epsilon}$), the overall sum must also be smaller in magnitude. This reasoning is incorrect.
    ANSW. ( 3.43 ) is a ratio not of the whole sum, but only of the positive contributions at different $\epsilon$ values. Each positive contribution decreases with increasing $\epsilon$, the same holds for their sum, so this ratio ( 3.43 ) also $\forall p_{max}<\infty$ is less than 1 and also (3.45 ) are completely justyfied. The same holds for negative contributions. Overall sum is not in question at this stage. The factor $p^{-\epsilon}$ is cited to give an idea. The computation takes in consideration correct contributions.
  • Analyzing the sum of positive terms in isolation from the sum of negative (like in (3.45 ) terms is not a valid operation
    ANSW Generally speaking I can agree but (see example two bullets above) $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$ tends exacly to the limit. So I do not understand the sentence: it“is not a valid operation”.
  • algebraic manipulations from ( 3.43 ) through (3.50) effectively treat the series as if it were a sum of positive terms, ignoring the crucial role of cancellation. This is a classic error when dealing with conditional convergence
@@
    ANSW In Lemma 4 it is used the function $Li(x)$, but, with whatsoever well behaved function the result of (3.19 ) is null because it depends from (3.17). So in Lemma 4 there is no reference to (PNTAP). The rationale of (3.20) is that of an algebraic trick to insert, with zero effective effect, the integral of PNT in (3.13). The difference in (3.20) means nothing because we are subtracting zero (i.e.(3.19 )) from (3.13). But it is useful in Lemma 6 to build not differences between primes distribution $\pi(x)$ and $Li(x)$ but ratios between functions involving $\pi(x)$ and $Li(x)$. In ratios asymptotic relations can be used at best. Of course not in differences. So I recommend a more careful reading and I think that the correct expectations of the reader will be perhaps satisfied.
@@
  • Computed results cannot be taken as proofs ANSW Yes I agree. Let us say they are a necessary condition not a sufficient one. However I think that the matching of the $t_{peak}$ of the peaks of (3.13) with the zeros of $L(s,\chi)$, and the matching of the floor with the computed one by (A.14 ) and (4.4) (see on figg, 3 through 13) is an interesting fact.
@@
  • The term “spectrum of primes” is not a rigorously defined concept ANSW Perhaps you are right. I borrowed in ([14]) this term by [16]. But anyhow a formula that gather zeros of L-functions and primes ( as (3.13) ), seems interesting.

### v16

@@ Appendix A Phase variations of $\xi(s)$, $\xi(s,\chi)$, $\zeta(s)$ and $L(s,\chi)$
With $\alpha_{1}=\alpha=0$, even $\chi_{primitive}$, or $\alpha_{1}=\alpha=1$, odd $\chi_{primitive}$. See (1.2 ). If we consider the factor (s1)(s-1) attached to $\zeta(s)$ like in (A.1), we can use (A.4) again putting $\alpha_{1}=0\ ,\ \alpha=2\ ,\ q=1$. As $z+1=\frac{s+\alpha}{2}\rightarrow z=\frac{2\epsilon+2\alpha-3}{4}+i\ \frac{t}{2}$, Stirling formula ( [8, p. 109,112]) allows us to write (A.4) for each of the 3 choices of $(\alpha,\alpha_{1})$:
@@
The $B_{j}$ are the Bernoulli numbers ( [8, p. 11] ), that vanish for odd $j$ while:
@@
where $\arg(z)$ is taken in the interval: $\pi<\arg(z)<\pi$ (see [8, p. 112].
@@
It is apparent that the difference between computation (3.16 ), upon substitution of (3.17) with (3.18) is absolutely convergent for sure from $\epsilon>0$ and conditionally convergent for $\epsilon=0$.
Proof. Let us look at (3.30), and, consider only one $h$-class primes till $p_{max}$ ( as in (3.21 ).
@@
If we consider all the $\phi(q)$ terms ( possible if $p_{max}<\infty$), even if we cannot conclude that the sum of all $\phi(q)$ contributions tend to zero ( as in Lemma 4), we can presume that the poor convergence of a single $h$-class primes is enanced when all $h$-classes are combined together. This seems confirmed by figg. 18, 19, and, 20.
@@ Appendix D Some feedback from readers
    ANSW.I agree on the necessity to fill this gap (not a mistake I think) in the logical path of the paper. In new version of section 3 I have shown that Euler product makes sense also in critical strip for primitive $L$-Functions. Equation (1.1) does not hold for $0<\Re(s)<1$, but, subsection 3.1 shows a way in order to exploit Euler product for variations, in particular phase variations, on small $t$-intervals. On the other hand ( even if it is not an argument) it seems to me worth to draw the attention on figg, 3 through 14 where it is apparent that (3.11), at least numerically, works rather well even with only $m=3$. In conclusion it seems to me that subsection 3.1 is a valid base for the developments of section 3. Anyhow I thank the reader for this observation that prompted me to delve deeper on subject.
@@
    ANSW. I agree with the statement, but it is not applicable.The “manipulation” of terms in (3.37 ) (or in others occasions (3.1), (3.21) and (3.23) ) is applied only at addends with index less than the running index. So on a set of finite numbers. Partial sums do not change (as it is stressed thoughout). To explain better. Consider an example $S_{j}=-1+1/2-1/3+1/4-.....=\sum_{j}(-1)^{j}/(j)\rightarrow-\ln(2)\ ,$ as $j\rightarrow\infty$.
    If we take the sum till $J_{max}$ and then we partition the addends with $j\leq J_{max}$ in positive $\sum^{+}_{J_{max}}$ and negatives $\sum^{-}_{J_{max}}$, then we can write $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$, and, we do not alter final result $\ln(2)$ for $J_{max}\rightarrow\infty$ transforming the limit in something like $\infty$-$\infty$. It is exactly the “manipulation” in expression : (3.37 ). While in (3.1), in (3.21), and, in (3.23), we manipulate to highlight some structure of the series, but , with indexes less than the running index . A completely rearranging of terms as in [1, p. 411] without change in final result is possible only where convergence is absolute. But, it is easy to see that there are infinite rearrangement of terms that preserve the sum $\ln(2)$ of example above. For example if we change odd with even terms, i.e. $S^{\prime}_{j}=-1/2+1-1/4+1/3-1/6+1/5-1/8+1/7-.....=-\ln(2)$.
@@
  • The argument in equations ( 3.46 ) through (3.53) rests on the idea that because each term in the series for $\epsilon>0$ is smaller in magnitude than the corresponding term for $\epsilon=0$ (due to the factor of $p^{-\epsilon}$), the overall sum must also be smaller in magnitude. This reasoning is incorrect.
    ANSW. ( 3.46 ) is a ratio not of the whole sum, but only of the positive contributions at different $\epsilon$ values. Each positive contribution decreases with increasing $\epsilon$, the same holds for their sum, so this ratio ( 3.46 ) also $\forall p_{max}<\infty$ is less than 1 and also (3.48 ) are completely justyfied. The same holds for negative contributions. Overall sum is not in question at this stage. The factor $p^{-\epsilon}$ is cited to give an idea. The computation takes in consideration correct contributions.
  • Analyzing the sum of positive terms in isolation from the sum of negative (like in (3.48 ) terms is not a valid operation
    ANSW Generally speaking I can agree but (see example two bullets above) $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$ tends exacly to the limit, provided the operation is carried on a finite terms set. So I do not understand the sentence: it“is not a valid operation”.
  • algebraic manipulations from ( 3.46 ) through (3.53) effectively treat the series as if it were a sum of positive terms, ignoring the crucial role of cancellation. This is a classic error when dealing with conditional convergence
@@
    ANSW In Lemma 4 it is used the function $Li(x)$, but, with whatsoever well behaved function the result of (3.22 ) is null because it depends from (3.20). So in Lemma 4 there is no reference to (PNTAP). The rationale of (3.23) is that of an algebraic trick to insert, with zero effective effect, the integral of PNT in (3.16). The difference in (3.23) means nothing because we are subtracting zero (i.e.(3.22 )) from (3.16). But it is useful in Lemma 6 to build not differences between primes distribution $\pi(x)$ and $Li(x)$ but ratios between functions involving $\pi(x)$ and $Li(x)$. In ratios asymptotic relations can be used at best. Of course not in differences. So I recommend a more careful reading and I think that the correct expectations of the reader will be perhaps satisfied.
@@
  • Computed results cannot be taken as proofs ANSW Yes I agree. Let us say they are a necessary condition not a sufficient one. In other words, for example, it makes sense to say that fig. 9 is compatible, i.e verifies, Lemma 6. Of course we cannot say it prove it. However I think that the matching of the $t_{peak}$ of the peaks of (3.16) with the zeros of $L(s,\chi)$, and the matching of the floor with the computed one by (A.14 ) and (4.4) (see on figg, 3 through 14) is an interesting fact.
@@
  • The term “spectrum of primes” is not a rigorously defined concept ANSW Perhaps you are right. I borrowed this term from [17], and stressed it in ([15]) . But anyhow a formula that gather zeros of L-functions and primes ( as (3.16) ), seems interesting.